how many gallons would a fish tank be if from all four sides were 20 inches and 10 inches
Answer by mananth(16946)

Volume = L*W*H
Volume = 20*20*10=4000 in^3
icu.in = 0.0043 gallon
4000*0.0043= 17.2 gallons